# EKF–SINDy: Empowering the extended Kalman filter with sparse identification of nonlinear dynamics

### v1

@@ 4 Conclusions
We have demonstrated this capacity by addressing two case studies. First, we addressed a linear non-autonomous system consisting of a shear building model, frequently used in vibration monitoring of civil structures, excited by real seismograms. Last, we faced a partially observed nonlinear system identifying in a first analysis the stiffness of the unobserved resonator, and in a second analysis the linear and quadratic coupling coefficients of the two resonators. Interestingly, we have showed how the procedure yields good outcomes also when the identification was operated outside the training range of SINDy. This demonstrates robustness and superior generalization capabilities of the proposed method with respect to alternative machine learning strategies, which rely on fully black-box, data-driven approaches, as, e.g., neural networks. We have also shown how the time delay embedding can be used to to uplift the dimensionality of the partial observations thus enabling the full description of the system dynamics.
Future and challenging directions involve applying the method to experimental data, which requires extending the approach to handle high-dimensional data. Given that SINDy is sensitive to data dimensionality, the strategy is to integrate the method with suitable dimensionality reduction techniques, such as, e.g., proper orthogonal decomposition and/or autoencoders champion2019data ; bakarji2023discovering . This integration aims to employ SINDy within a reduced space while preserving the parametric dependency as in conti2023reduced . To further enhance the use of the procedure, we will also consider to employ automatic tuning procedures based on genetic algorithms art:Rapp03 or on swarm intelligence art:Laamari15 to perform the automatic tuning of the filter.
@@
https://github.com/ContiPaolo/EKF-SINDy EKFSINDY_repo .
@@ Acknowledgments
We now report further information about the tuning of the EKF for the considered cases of study. In all cases, we have initialised the state covariance matrix $\mathbf{P}(0)$, the process noise covariance matrix $\mathbf{Q}$, and the observation noise covariance matrix $\mathbf{R}$ as diagonal matrices. In Tabs. A.1-A.3, the diagonal terms of these matrices are gathered.
@@
Table A. 1: Initial covariance $\mathbf{P}(0)$, diagonal terms. In the left column, we specify the augmented state quantities to which these values are related.
       Shear building $x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}$ $5\cdot 10^{-3}$ $k_{1}$ $10^{-8}$ Nonlinear dynamic system $\bar{k}_{2}=1.2$ $x_{1},x_{2},x_{3},x_{4}$ $10^{-7}$ $k_{2}$ $10^{-7}$ $\bar{k}_{2}=2.3$ $x_{1}$ $10^{-10}$ $x_{2}$ $10^{-10}$ $x_{3}$ $10^{-8}$ $x_{4}$ $10^{-8}$ $k_{2}$ $2\cdot 10^{-6}$ $\alpha$, $\beta$ $x_{1}$ $5\cdot 10^{-10}$ $x_{2}$ $5\cdot 10^{-10}$ $x_{3}$ $2\cdot 10^{-8}$ $x_{4}$ $3\cdot 10^{-8}$ $\alpha$ $3\cdot 10^{-1}$ $\beta$ $3\cdot 10^{-1}$ Table A. 2: Process noise $\mathbf{Q}$, diagonal terms. In the left column, we specify the augmented state quantities to which these values are related.        Shear building $x_{1}$ $10^{-3}$ $x_{2}$ $10^{-3}$ $\dot{x}_{1}$ $10^{-3}$ $\dot{x}_{2}$ $10^{-3}$ $\ddot{x}_{1}$ $10^{3}$ $\ddot{x}_{2}$ $10^{3}$ Nonlinear dynamic system $\bar{k}_{2}=1.2$ $z_{1}$ $10^{-7}$ $\bar{k}_{2}=2.3$ $z_{1}$ $5\cdot 10^{-4}$ $\alpha$, $\beta$ $z_{1}$ $3\cdot 10^{-1}$ Table A. 3: Measurement noise $\mathbf{R}$, diagonal terms. In the left column, we specify the observed quantities to which these values are related.

### v2

@@ 4 Conclusions
We have demonstrated this capacity by addressing two case studies. First, we addressed a linear non-autonomous system consisting of a shear building model, frequently used in vibration monitoring of civil structures, excited by real seismograms. Last, we faced a partially observed nonlinear system identifying in a first analysis the stiffness of the unobserved resonator, and in a second analysis the linear and quadratic coupling coefficients of the two resonators. Interestingly, we have showed how the procedure yields good outcomes also when the identification was operated outside the training range of SINDy. This demonstrates robustness and superior generalization capabilities of the proposed method with respect to alternative machine learning strategies, which rely on fully black-box, data-driven approaches, as, e.g., neural networks. We have also shown how the time delay embedding can be used to uplift the dimensionality of the partial observations thus enabling the full description of the system dynamics.
Future and challenging directions involve applying the method to experimental data, which requires extending the approach to handle high-dimensional data. Given that SINDy is sensitive to data dimensionality, the strategy is to integrate the method with suitable dimensionality reduction techniques, such as, e.g., proper orthogonal decomposition and/or autoencoders [13, 11]. This integration aims to employ SINDy within a reduced space while preserving the parametric dependency as in [14]. To further enhance the use of the procedure, we will also consider to employ automatic tuning procedures based on genetic algorithms [59] or on swarm intelligence [60] to perform the automatic tuning of the filter.
@@
https://github.com/ContiPaolo/EKF-SINDy [28].
@@ Acknowledgments
Sparsity promoting parameter
The value of the threshold parameter $L$ affects the sparsity of the discovered functional relations. Conversely, it has not been necessary to adjust the regularisation strength $\delta_{r}$, which has been fixed at the default value of $0.05$ in the PySINDy package. In Tabs. A.1, we report the value of $L$ used for each case study.
Shear building Nonlinear system ($k_{2}$) Nonlinear system ($\alpha$ and $\beta$)
$10^{2}$ $5\cdot 10^{-4}$ $10^{-3}$
Table A. 1: Sparsity promoting threshold parameter $L$.
The resulting functional relations are detailed in the following, specifying the order of the polynomial terms (linear, quadratic or cubic) employed by the trained SINDy models, and the state variables relevant to each case study.
Shear building under seismic excitations
The state vector $\boldsymbol{\varkappa}$ collects the floor displacements, velocities, and interstorey stiffness $k$, namely $\boldsymbol{\varkappa}=[x_{1},x_{2},\dot{x}_{1},\dot{x}_{2},k]^{\top}$. In this way, SINDy can be directly applied to model the evolution of the state vector $\boldsymbol{\varkappa}$ as a first order differential equation, where state variables have been suitably rescaled to have comparable magnitude. Polynomial terms up to the second order are considered for the state variables, while extra linear terms are used to account for the external forcing that makes the system non-autonomous. Specifically, the forcing terms $b_{1}$ and $b_{2}$ are applied respectively to the first and second interstoreys.
The identified SINDy model is reported here at the second order consistently with the form of Eq. (22) ruling the dynamics of the system. For the first order formulation, the reader can refer to the GitHub repository [28].
$\begin{split}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ddot{x}_{1}=\\ -0.716x_{1}+0.446x_{2}-2.25\dot{x}_{1}+1.112\dot{x}_{2}&\qquad\text{(linear)}\\ +308259736.25b_{1}-308259740.5b_{2}&\qquad\text{(forcing)}\\ -0.088{x_{1}}^{2}+0.055x_{1}x_{2}+2.984x_{1}\dot{x}_{1}-1.883x_{1}\dot{x}_{2}-351.04x_{1}k-1.844x_{2}\dot{x}_{1}&\qquad\text{(quadratic)}\\ +1.164x_{2}\dot{x}_{2}+174.107x_{2}k-0.534{\dot{x}_{1}}^{2}+0.598\dot{x}_{1}\dot{x}_{2}-3.768\dot{x}_{1}k-0.164{\dot{x}_{2}}^{2}+1.776\dot{x}_{2}k&\qquad\text{(quadratic)}\end{split}$ (A.1a)
$\begin{split}\ddot{x}_{2}=\\ -0.557x_{1}+0.346x_{2}+0.922\dot{x}_{1}-1.023\dot{x}_{2}&\qquad\text{(linear)}\\ +307191775.013b_{1}-307191779.264b_{2}&\qquad\text{(forcing)}\\ -3.33{x_{1}}^{2}+4.031x_{1}x_{2}-0.01x_{1}\dot{x}_{1}+193.088x_{1}k-1.218{x_{2}}^{2}+0.056x_{2}\dot{x}_{1}&\qquad\text{(quadratic)}\\ -0.03x_{2}\dot{x}_{2}-188.664x_{2}k-0.09{\dot{x}_{1}}^{2}+0.036\dot{x}_{1}\dot{x}_{2}+3.186\dot{x}_{1}k+0.014{\dot{x}_{2}}^{2}-2.861\dot{x}_{2}k&\qquad\text{(quadratic)}\end{split}$ (A.1b)
Nonlinear dynamical system, estimation of the stiffness $k_{2}$ of the hidden oscillator
The state vector $\boldsymbol{\varkappa}$ collects the time delay coordinates and the $k_{2}$, namely $\boldsymbol{\varkappa}=[x_{1},x_{2},x_{3},x_{4},k_{2}]^{\top}$. Considering a polynomial terms up to the third order, the total number of library functions is $55$. According to the set threshold level $L$, SINDy employs: $4$ terms to model $\dot{x}_{1}$; $5$ terms for $\dot{x}_{2}$; $7$ terms for $\dot{x}_{3}$; $9$ terms for $\dot{x}_{4}$.
$\begin{split}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\dot{x}_{1}=\\ -0.007x_{1}-1.027x_{2}-0.001x_{3}-0.009x_{4}&\qquad\text{(linear)}\end{split}$ (A.2a)
$\begin{split}\dot{x}_{2}=\\ 1.004x_{1}-0.020x_{2}-0.150x_{3}&\qquad\text{(linear)}\\ -0.001x_{1}k_{2}+0.004x_{3}k_{2}&\qquad\text{(quadratic)}\end{split}$ (A.2b)
$\begin{split}\dot{x}_{3}=\\ 0.005x_{1}+0.157x_{2}-0.002x_{3}-1.246x_{4}&\qquad\text{(linear)}\\ -0.001x_{1}k_{2}-0.006x_{2}k_{2}+0.019x_{4}k_{2}&\qquad\text{(quadratic)}\end{split}$ (A.2c)
$\begin{split}\dot{x}_{4}=\\ 0.262x_{1}-0.049x_{2}-0.061x_{3}-0.011x_{4}&\qquad\text{(linear)}\\ -0.189x_{1}k_{2}+0.044x_{2}k_{2}+0.834x_{3}k_{2}&\qquad\text{(quadratic)}\\ -0.001x_{2}k_{2}^{2}+0.002x_{3}k_{2}^{2}&\qquad\text{(cubic)}\end{split}$ (A.2d)
Nonlinear dynamical system, estimation of the coupling parameters $\alpha$ and $\beta$
The state vector $\boldsymbol{\varkappa}$ collects the time delay coordinates and the coupling parameters, namely $\boldsymbol{\varkappa}=[x_{1},x_{2},x_{3},x_{4},\alpha,\beta]^{\top}$. Considering a polynomial terms up to the third order, the total number of library functions is $83$. According to the set threshold level $L$, SINDy employs: $9$ terms to model $\dot{x}_{1}$; $12$ terms for $\dot{x}_{2}$; $14$ terms for $\dot{x}_{3}$; $29$ terms for $\dot{x}_{4}$.
$\begin{split}\dot{x}_{1}=\\ -0.006x_{1}-1.013x_{2}-0.001x_{3}-0.034x_{4}&\qquad\text{(linear)}\\ +0.001x_{4}\beta&\qquad\text{(quadratic)}\\ -0.001x_{2}\alpha^{2}-0.001x_{2}\alpha\beta+0.001x_{4}\alpha^{2}+0.001x_{4}\alpha\beta&\qquad\text{(cubic)}\end{split}$ (A.3a)
$\begin{split}\dot{x}_{2}=\\ +0.996x_{1}-0.014x_{2}-0.343x_{3}&\qquad\text{(linear)}\\ +0.002x_{1}\alpha+0.002x_{3}\alpha+0.005x_{3}\beta&\qquad\text{(quadratic)}\\ -0.028x_{1}\alpha^{2}+-0.007x_{1}\alpha\beta-0.004x_{1}\beta^{2}+0.007x_{3}\alpha^{2}+0.021x_{3}\alpha\beta+0.011x_{3}\beta^{2}&\qquad\text{(cubic)}\end{split}$ (A.3b)
$\begin{split}\dot{x}_{3}=\\ -0.003x_{1}+0.338x_{2}-0.001x_{3}-1.924x_{4}&\qquad\text{(linear)}\\ +0.002x_{1}\alpha-0.002x_{2}\beta+0.011x_{4}\alpha+0.015x_{4}\beta&\qquad\text{(quadratic)}\\ -0.042x_{2}\alpha^{2}-0.021x_{2}\alpha\beta-0.012x_{2}\beta^{2}+0.029x_{4}\alpha^{2}+0.059x_{4}\alpha\beta+0.030x_{4}\beta^{2}&\qquad\text{(cubic)}\end{split}$ (A.3c)
$\begin{split}\dot{x}_{4}=\\ +0.114x_{1}-0.004x_{2}+1.849x_{3}+0.029x_{4}&\qquad\text{(linear)}\\ +8.899x_{1}^{2}-0.422x_{1}x_{4}+0.089x_{1}\alpha+0.035x_{1}\beta+0.290x_{2}x_{3}-0.008x_{2}\alpha&\qquad\text{(quadratic)}\\ -0.478x_{3}\alpha-0.285x_{3}\beta+0.916x_{4}^{2}+0.497x_{4}\alpha-0.003\alpha^{2}-0.001\beta^{2}&\qquad\text{(quadratic)}\\ -1.242x_{1}\alpha^{2}-0.244x_{1}\alpha\beta-0.230x_{1}\beta^{2}+0.086x_{2}\alpha^{2}-0.202x_{2}\alpha\beta&\qquad\text{(cubic)}\\ -0.759x_{3}\alpha^{2}-0.460x_{3}\alpha\beta+1.501x_{3}\beta^{2}+0.950x_{4}\alpha^{2}&\qquad\text{(cubic)}\\ -0.392x_{4}\alpha\beta-0.006\alpha^{3}+0.018\alpha^{2}\beta+0.002\beta^{3}&\qquad\text{(cubic)}\end{split}$ (A.3d)
APPENDIX B
@@
We report further information about the tuning of the EKF for the considered cases of study. In all cases, we have initialised the state covariance matrix $\mathbf{P}(0)$, the process noise covariance matrix $\mathbf{Q}$, and the observation noise covariance matrix $\mathbf{R}$ as diagonal matrices. In Tabs. B.1-B.3, the diagonal terms of these matrices are gathered.
@@
Table B. 2: Initial covariance $\mathbf{P}(0)$, diagonal terms. In the left column, we specify the augmented state quantities to which these values are related.
       Shear building $x_{1},x_{2},\dot{x}_{1},\dot{x}_{2}$ $10^{-4}$ $k$ $10^{-8}$ Nonlinear dynamic system $\bar{k}_{2}=1.2$ $x_{1},x_{2},x_{3},x_{4}$ $10^{-7}$ $k_{2}$ $10^{-7}$ $\bar{k}_{2}=2.3$ $x_{1}$ $10^{-10}$ $x_{2}$ $10^{-10}$ $x_{3}$ $10^{-8}$ $x_{4}$ $10^{-8}$ $k_{2}$ $2\cdot 10^{-6}$ $\alpha$, $\beta$ $x_{1}$ $5\cdot 10^{-10}$ $x_{2}$ $5\cdot 10^{-10}$ $x_{3}$ $8\cdot 10^{-8}$ $x_{4}$ $8\cdot 10^{-8}$ $\alpha$ $10^{-11}$ $\beta$ $10^{-11}$ Table B. 3: Process noise $\mathbf{Q}$, diagonal terms. In the left column, we specify the augmented state quantities to which these values are related.        Shear building $x_{1}$ $5\cdot 10^{-1}$ $x_{2}$ $5\cdot 10^{-1}$ $\dot{x}_{1}$ $5\cdot 10^{-3}$ $\dot{x}_{2}$ $5\cdot 10^{-3}$ $\ddot{x}_{1}$ $5$ $\ddot{x}_{2}$ $5$ Nonlinear dynamic system $\bar{k}_{2}=1.2$ $z_{1}$ $10^{-7}$ $\bar{k}_{2}=2.3$ $z_{1}$ $5\cdot 10^{-4}$ $\alpha$, $\beta$ $z_{1}$ $3\cdot 10^{-1}$ Table B. 4: Measurement noise $\mathbf{R}$, diagonal terms. In the left column, we specify the observed quantities to which these values are related.
As mentioned, Kalman filter tuning has been operated through a trial-and-error procedure. Some practical guidelines can be used to expedite this procedure. These guidelines are derived by experience, and account for the probabilistic meaning of the process $\mathbf{Q}$ and observation $\mathbf{R}$ noise covariance matrices. The following heuristic discussion must thus be intended more as a practical guidance for the reader, than as a systematic approach to Kalman filter tuning.
First, it is advisable to set the entries of $\mathbf{R}$ based on the specifications of the acquisition system. Next, the identification outcome should be checked for a random initialisation of $\mathbf{Q}$. If the filter is unable to update the mean of the model parameter predictions $\hat{\boldsymbol{\phi}}$ during the acquisition window, the entries of $\mathbf{R}$ are probably too high. If the value of $\hat{\boldsymbol{\phi}}$ is instead sensitive to the acquisitions, attention should then be directed to $\mathbf{Q}$. Examining again the identification outcomes, if the confidence intervals do not reduce during the analysis, it is likely that the entries of $\mathbf{Q}$ are set too high, causing the filter to overly trust measurements and neglect model predictions. This precludes the filter from gaining confidence in its predictions, and thus reducing the confidence intervals. Therefore, it is advisable to lower the process noise parameters. In doing that, one should initially operate large changes in the entries of $\mathbf{Q}$ (even by an order of magnitude), and perform a fine tuning in the later trials. However, excessive reduction of the process noise parameters may cause divergence in the filter estimates.